. Adam Smith is a portfolio manager with Point72 Investments, a U.S.-based asset management firm. Smith is considering using options to enhance portfolio returns and control risk. He asks his junior analyst, Tommy Hilfiger, to help him. Hilfiger collected the current market prices and data of selected instruments related to Lotus stock in Table 2. According to Table 2, which of the following should be Smith's arbitrage strategy? (Choose the best answer)

Table 2

European call option on Lotus equity

$3

European put option on Lotus equity

$3

Lotus equity price

$50

Time to expiration of options

3 months

Exercise price of options

$50

Risk-free rate

10%

Dividend to be paid in 3 months

$3.32

"

Long put option, long stock, short call option, and lending money

" short put option, short stock, short call option, and lending money"

" long put option, long stock, short call option, and borrowing money"

" short the put option, short the stock, long the call option, and lending money"

" short put option, long stock, long call option, and lend money"

" short put option, short stock, long call option, and borrow money"
Answer #1

Long put option, long stock, short call option, and lending money

Using the put-call parity
S+P=C+X/(1+r)^t+Present Value of Dividends

LHS=S+P=50+3=53

RHS=C+X/(1+r)^t+Present Value of Dividends=3+50/1.1^(3/12)+3.32/(1+10%)^(3/12)=55.06453206

As LHS<RHS, buy LHS and sell RHS